OpenStudy (anonymous):

Use elimination to solve the system 3x - 5y = 8 and 4x - 7y = 10. What is the value of x?

OpenStudy (anonymous):

3x-5y=8 *4 4x-7y=10 *3 12x-20y=32 12x-21y=30 ----------- - y=2 3x-5(2)=8 3x=18 x=6